7, A ball A is thrown upward at the same time with a ball B. The initial speed of A is half that of B. What is the distance of A from the top of B?


1
Expert's answer
2021-09-13T00:04:47-0400

"\text{We know that $v^2 =u^2+2gh$, at maximum height velocity is 0, therefore we have}\\
U_A^2 = -2gh_A-(1)\\
U_B^2 = -2gh_B\\
\text{But $U_A=\frac{U_B}{2}$}\\
\implies 4U_A^2 = -2gh_B\\
h_B= -\frac{4U_A^2}{2g}\\
\text{From (1), we have that $h_A=-\frac{U_A^2}{2g}$}\\
\therefore h_B = 4h_A\\
\text{Hence, the difference between the distance of A and B is 4A-A= 3A }"